OpenStudy (anonymous):

Leo bought a bulldozer for $63,103. The value of the bulldozer depreciated at a constant rate per year. The table below shows the value of the bulldozer after the first and second years:

OpenStudy (anonymous):

Year 1 2 Value (in dollars) 58,054.76 53,410.38

OpenStudy (anonymous):

Which function best represents the value of the bulldozer after t years? f(t) = 58,054.76(0.92)t f(t) = 63,103(0.08)t f(t) = 63,103(0.92)t f(t) = 58,054.76(0.08)t

OpenStudy (anonymous):

the table looks like this : year 1 2 valued money 58,054.76 53,410.38

OpenStudy (anonymous):

@Green_Plan @ganeshie8 @random231 @radar

random231 (random231):

the first one is the answer because the rate of depreciation is 53,410.38/58,054.76 = 0.92 so f(t)= 58,054.76*0.92*t

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (anonymous):

thats what i thought too.

OpenStudy (anonymous):

One more question, why wouldnt it be 63,103?

OpenStudy (anonymous):

@random231 ?

OpenStudy (anonymous):

ok, just checking.

random231 (random231):

just a second.. @naturegirl

random231 (random231):

im sorry i said it wrong i just saw the table!!!!!!! :((((((((sorry

random231 (random231):

itll be the third one... to be purfect

random231 (random231):

but i must add the first one is not entirely wrong!!!

OpenStudy (anonymous):

Ok, lol thank you for all your help!

random231 (random231):

im sorry!!! :(

random231 (random231):

for misguiding u!